OpenStudy (anonymous):

how to graph x+4y=16

OpenStudy (anonymous):

solve for y 4y=16-x y=-(1/4)x+4 This is your standard y=mx+b form x-intercept is at x=4 the slope is (-1/4) (the graph does downward)

OpenStudy (anonymous):

a very easy procedure is to fine intercepts for x-intercept put y=0 and solve for x similarly for y-intercept u got two points (a,0) and (0,b) u r done put that on graph and join them

OpenStudy (anonymous):